Hi!
Could you please explain how do I do this probblem.
a) find the center-radius from of the equation of each circle
b) graph it
center (5-4) radius 7
How do I know which way the line is going to be?
Thanks

Answer by Nate(3500)

Your question:
Could you please explain how do I do this problem.
a) find the center-radius from the equation of each circle
b) graph it
center (5-4) radius 7
How do I know which way the line is going to be?
Thanks
--------------------------------------------------------
My answer:
The basic formula for a circle is this: (x-h)^2+(y-k)^2=r^2
"h" is the x value from the center point and "k" is the y value from the center point.........(ex.) the ordered pair for the center point is (h,k)
If your center is (5,4) and the radius is 7, then your equation should look like (x-5)^2+(y-4)^2=49
To graph this, you would graph the center. Then, you would go up, down, right, and to the left of that point for 7 units. If you had a compass, you would just connect the points with a better circle look. If not, then good luck freehanded. The line would be round hence that it is a cirlce.
N8